distance = \sqrt{(North position)^{2} + (East position)^{2}}

Parable  Special Relativity
North measured in miles  Time measured in seconds
East measured in feet  Distance measured in meters
Conversion k = 5280 feet/mile  Conversion c = 2.99792458 × 10^{8} m/s
N(feet) = kN(miles)  t(m) = ct(s)
E(miles) = N(feet)/k  t(s) = t(m)/c
North interval ΔN = N_{2} − N_{1}  Space interval Δx = x_{2} − x_{1}
East interval ΔE = E_{2} − E_{1}  Time interval Δt = t_{2} − t_{1}
(Displacement Interval)^{2} = (kΔN)^{2} + (ΔE)^{2}  (SpaceTime Interval)^{2} = (cΔt)^{2} − (Δx)^{2}
(Displacement Interval)^{2} = (ΔN)^{2} + (ΔE)^{2}  (SpaceTime Interval)^{2} = (Δt)^{2} − (Δx)^{2}

(SpaceTime Interval)^{2} = (cΔt)^{2} − (Δx)^{2}

(SpaceTime Interval)^{2} = (Δt)^{2} − (Δx)^{2}

L = 4.07 × 10^{16} meters
(SpaceTime Interval)^{2}  = (Δt)^{2} − (Δx)^{2}
(SpaceTime Interval)^{2} + (Δx)^{2} = (Δt)^{2}
(SpaceTime Interval)^{2} + 0  = (Δt)^{2}
(SpaceTime Interval)^{2}  = (Δt)^{2}
SpaceTime Interval  = Δt
c = \frac{1 meter of space}{1 meter of time} = 1

c = \frac{1 year of space}{1 year of time} = 1